# Exploring the Performance-Reproducibility Trade-off in Quality-Diversity

### v1

@@ VI-A Results on Existing UQD Benchmarks
Figure 6: Reproducibility-maximisation benchmark results: in these tasks, the fitness is always 00, so we only display the reproducibility averaged over the archive. We report the median and quartiles across $10$ seeds.
@@ VI-C Results on Extended UQD Benchmarks
Figure 8: Trade-off benchmark results: we report the fitness averaged over the archive corresponding to the performance-reproducibility profiles in Figure 7. The green-coloured areas indicate fitness values that algorithms are expected to converge to. We report the median and quartiles across $10$ seeds.
@@ VII Conclusion and Discussion
Acknowledgements
We would like to thank all the other members of the Adaptive and Intelligent Robotics Lab for their useful feedback.

### v2

@@ VI-A Results on Existing UQD Benchmarks
Figure 6: Reproducibility-maximisation benchmark results: in these tasks, the fitness is always 00, so we only display the reproducibility averaged over the archive. For this metric, a higher score is better, and the maximum value is $1$. The vertical lines show the median across $10$ replications.
@@ VI-C3 Avoidable and Unavoidable Peak
Figure 8: Trade-off benchmark results: we report the fitness averaged over the archive corresponding to the performance-reproducibility profiles in Figure 7. The green-coloured areas indicate fitness values that algorithms are expected to converge to. The vertical lines show the median across $10$ replications.